using the vowels of the english alphabet, how many different possible 3-letter code words can be formed?
Answer by jim_thompson5910(35256)

There are 5 vowels: a, e, i, o, u
I'm excluding y.


Assuming repeats are allowed, you can do 5*5*5 = 5^3 = 125 different code words.


Assuming repeats are NOT allowed, you can do 5*4*3 = 60 different code words.


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim